Hi all,
I am sorry for my question, but I don't know how solve some equations.
I need isolate "s" in this problem:
P=f/2(1+sqrt(1+4s^2/f2))
Thanks in advance!


Answer by stanbon(75887)

I need isolate "s" in this problem:
P=f/2(1+sqrt(1+4s^2/f^2))
----
2 = f[1+sqrt((f^2+4s^2)/f^2)]
-----
2 = f[1+(1/f)sqrt(f^2+4s^2)]
====
2 = [f+sqrt(f^2+4s^2)]
-----
2-f = sqrt(f^2+4s^2)
----
4-4f+f^2 = f^2 + 4s^2
----
4s^2 = 4-4f
-----
s^2 = 1 - f
-----
s = sqrt(1-f)
------------------
Cheers,
Stan H.
-----------